For this reaction

2 H2O (g) −⇀ 2 H2 (g) + O2 (g)

the equilibrium concentrations were found to be [H2O] = 0.250 M, [H2] = 0.410 M,

and [O2] = 0.750 M.

What is the equilibrium constant for this reaction?


Kc = ?


1
Expert's answer
2021-06-23T05:17:40-0400

Kc = "\frac{[H_2O]^2}{[H_2]^2 [O_2]}" = "\frac{[0.250]^2]}{[0.410]^2 [0.750]}"


Kc = "\frac{0.0625}{0.1681×0.750}" = "\frac{0.0625}{0.126075}" = 0.495736


Kc = 0.496 approx